5x + 5y - 4z = 13
5x + 3y + z = - 3
3x - 2y - z = - 10
x=___
y=___
z=___
solve the system

Answer by Alan3354(69443)

5x + 5y - 4z = 13
5x + 3y + z = -3
3x - 2y - z = -10
----------
Add equations 2 & 3
---
Multiply eqn 2 by 4, then add eqn 1.
========
z is eliminated.
******************************
Do not enter this.
x=___
y=___
z=___
We have no use for it.